Spouse Details - Michelle Hart

            Corey’s wife, Michelle, quit her full-time job with Queensland Health one year ago and enrolled full-time in a Bachelor of International Studies degree at Griffith University. Michelle also has a small part-time job working ten hours per week at a local Toowong coffee shop.

            Michelle’s assessable income for the 2018 income year consisted of gross wages totalling $12,180 (this amount includes PAYG withholding tax of $1,420 deducted) and interest income of $26. Michelle also received a fully franked dividend of $490 from Wesfarmers Ltd during the 2018 income year.

            Michelle has no other assessable or exempt income, no reportable fringe benefits, and no reportable employer superannuation contributions.

            Michelle incurred allowable deductions of $278 in relation to her salary and wage income.

Please identify what is tax deductible for Michelle.
Answer #1

Taxable income:12180+26-278=11928

tax on above income=(9700*10%)+((11928-9700)*12%)=1237

Tax already deducted:1420

Dividend shall not be considered under total income since it is exempt from tax.

No tax is deductible from salary of Michelle.